40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 227 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  5450   Sun Sep 18 15:57:00 2011 KojiUpdateIOOThe PZT driver engaged to PZT1

[Koji Kiwamu]

The pzt driver for PZT1 has been installed.
As there was unknown resistive connection in the vacuum chamber as described below,
the PZT out cable at the PJ driver module should always be disconnected.
The sensor cables have no problem to be connected to the controller.
In fact, they are a good monitor for the state of the PZTs.

In this configuration, Pitch and Yaw direction of PZT1 is under the control of the EPICS value as we expected.


Details:

- At the beginning, we tested the PZT driver output with low voltage level (~10V). We did not see any oscillation of the opamps.
  The pitch output was observed to be OK, while the YAW output exhibited a half of the expected output voltage.
  The opamp was holding correct voltage, however the voltage after the 1K output resister was about a half.
  This indicated there was a voltage division happening.

- The cause of the voltage division was tracked. We found that the yaw red (=hot) line is connected to pitch black
  in the vacuum chamber with a resistance of 1.4kOhm. The black cables are shorted to the ground level in the PJ driver.

- We decided to unplug the PJ's cable so that we can isolate the black cables while hoping the PZTs were drived only
  by the red and white cables. And they did.

- This means that we should not connect the PZT driving cable to the PJ's driver. The sensors have no problem to be connected.

- Pinouts:

DSUB25
|. .|
|. .|
|. o|  5
|o  | 17
|  o|  4 
|o  | 16   Yaw Black
|  o|  3   Pitch Black
|o  | 15   Yaw White
|  o|  2   Yaw Red
|o  | 14   Pitch White
 \ o|  1   Pitch Red
  \-+

* Pitch White and Yaw White are connected to the ground at the amplifier side.
* Yaw Red and Pitch Black is connected with 1.4kOhm and isolated from the others.


  5449   Sun Sep 18 15:34:09 2011 KojiUpdateAdaptive FilteringModifications to LSC, RFM models, added OAF model

[Koji Kiwamu]

This modification of the LSC model made the rows of the LSC output matrix shifted. This caused the ASS scripts nonfunctional.

Kiwamu fixed the channel names in the ASS script.

Quote:

[Jenne, Mirko, with supervision from Jamie]

I modified the c1lsc model to have shmem outputs that go from the degrees of freedom to the OAF, and shmem inputs from the OAF's output to sum into the DoFs, just like Yoichi's FF stuff.  I also removed the old OAF_OUT, because it would only allow me to select one DoF at a time, and I will eventually want the ability to do multiple amounts of OAFing at the same time.  Hopefully.

 

  5448   Sun Sep 18 14:08:52 2011 ranaUpdateSUSCalibration plan for the oplevs

We don't need a high quality calibration for the optical levers. ~50% accuracy is fine.

For that you can use the OSEM calibration of ~1.7 V/mm (its less than 2 since the OSEMs have been degrading) or you can use the cavity power method that Kakeru used; it worked just fine. There's no benefit in trying for a 1% number for optical levers.

  5447   Sat Sep 17 14:04:45 2011 KojiUpdateIOOPZT1 driver in place

The PZT driver is now in place. The actual PZTs are not connected yet!

It is accommodated on Ben's connector adapter board.

The panel has additional connectors now: two inputs and a power supply connector.

The supply voltage is +/-30V (actual maximum +/-40V), and the input range is +/-10V
which yields the output range of -5V to 30V. The gain of the amplifier is +2.

It is confirmed that the HV outputs react to the epics sliders although the PZT connector is not connected yet
so as not to disturb the locking activity.

When we engage the PZT connector, we should check the HV outputs with an oscilloscope to confirm they
have no oscillation with the capacitances of the PZTs together with the long cable.

Attachment 1: P9171579.JPG
P9171579.JPG
Attachment 2: P9171580.JPG
P9171580.JPG
  5446   Sat Sep 17 02:07:10 2011 kiwamuUpdateSUSETMX input matrix : bad YAW-SIDE coupling

Excited all the optics. They will be automatically back after 5 hours.

Sat Sep 17 02:02:07 PDT 2011
1000285342

Quote from #5444

It needs one more kick and free swinging test.

 

  5445   Sat Sep 17 01:53:41 2011 KeikoUpdateLSCPOY and POP beams clipped

 Keiko, Paul, Kiwamu

We found that POP beam is clipped by the steering mirrors inside the tank. POY beam is also likely to be clipped inside. Also the hight of POY beam is too high (about 5 cm higher than the normal paths) at the first lens. These imply the input pointing is bad.

  5444   Fri Sep 16 23:22:36 2011 kiwamuUpdateSUSETMX input matrix : bad YAW-SIDE coupling

With the new input matrix, it looks like YAW and SIDE are not quite decoupled on ETMX.

It needs one more kick and free swinging test.

 

- - - details

 To see what exactly is going on, I changed the input matrix from the default to the new one, which Jenne computed (#5421) on ETMX.

I started putting the elements of the input matrix from POS through SIDE, one by one.

It seemed that POS and PIT worked fine. However the YAW signal looks containing a lot of the SIDE signal.

Similar to YAW, SIDE also interact with the YAW motion and somehow rings up both YAW and SIDE signals as Jenne reported ( #5438).

So right now the YAW and SIDE rows are partially reburted to the default elements in order to avoid ringing up.

Quote from #5438

but ETMX and BS were not good at all.  ETMX was ringing up when I turned on the damping. 

 

  5443   Fri Sep 16 22:51:52 2011 PaulUpdateSUSCalibration plan for the oplevs

 In order to estimate the amount of noise that the oplevs are injecting into the GW channel, we first need to calibrate oplev signals in terms of angular change in the optic. I said in my previous post that there wasn't a calibration factor for OSEM values to radians, but I found that Kakeru had estimated this in 2009 - see entry 1413. However, Kakeru found that this was quite a rough estimate, and that it didn't agree with his calibrated oplev values well. He does quote the 2V/mm calibration factor for the OSEM readings though - does anyone know the provenance of this factor? I searched for OSEM calibration and found nothing.

 
Kiwamu and Suresh suggested a way to calibrate the oplevs without needing to calibrate the OSEMs in the way that Kakeru describes in entry 1413. This should give a calibration for the OSEMs _and_ the oplevs in fact. The method should be as follows:
 
1) Change the coil driver values in DC to give tip or tilt the optic. Measure the resulting change in spot position at a known distance from the optic, perhaps just using a ruler. Record the spot position and OSEM values for each coil driver value. This will definitely require a smaller spot size, so I'll implement the new telescopes first.
 
2) Knowing the length of the lever arm from the optic to the spot measurement position, we can calibrate the OSEM values to radians.
 
3) We can now put the beam onto the oplev QPD, and either change the coil driver values again in the same way (but over a smaller range), or excite the test mass in pitch or yaw, this time measuring both the OSEM values and the oplev QPD values. Since we can already convert from OSEM values to radians, we can now convert from oplev values to radians too.
 
4) I should be careful to consider the input sensing matrix for both the OSEMs and the oplevs in these measurements. Should I divide those out of the calibration to avoid that if they change the calibration factor changes too?
  5442   Fri Sep 16 22:11:21 2011 PaulUpdateSUSITMY transfer function

First of all I moved the lenses on the ITMY/SRM oplev path to get a smaller spot size on the QPDs. I couldn't get the beam analyzer to work though, so I don't know quite how successful this was. The software brought up the error "unable to connect to framegrabber" or something similar. I don't think the signal from the head was being read by the software. I will try to get the beam analyzer working soon so that we can characterize the other oplev lasers and get decent spot sizes on the QPDs. I searched the elog for posts about the analyzer, and found that it has been used recently, so maybe I'm just doing something wrong in using it. 

After this I measured the transfer function for the ITMY oplev yaw. I did a swept sine excitation of the ITMY in yaw with an amplitude of 500, and recorded the OSEM yaw values and the oplev yaw values. This should show a flat response, as both the QPD and the OSEMS should have flat frequency response in the measurement band. This measurement should therefore just yield a calibration from OSEM yaw to oplev yaw. If the OSEM yaw values were already calibrated for radians, we would then immediately have a calibration from oplev yaw values to radians. However, as far as I'm aware, there is not a calibration factor available from OSEM yaw values to radians. Anyway, the TF I measured did not appear to be very flat (see attached plot). Kiwamu suggested I should check the correlation between the OSEM measurements and the oplev QPD measurements - if the correlation is less than 1 the TF is not reliable. Indeed the coherence was poor for this measurement. This was probably because at frequencies above the pendulum frequency, the excitation amplitude of 500 was not enough to cause a measurable change in the optic angle. So, the plot attached is not very useful yet, but I learned something while making it.

 

Attachment 1: ITMY_osem_to_oplev_TF.pdf
ITMY_osem_to_oplev_TF.pdf
  5441   Fri Sep 16 21:36:25 2011 KeikoUpdateLSCPOY11 and POY55 were added

 New channels, POP55 and POY11 are connected to the rack and now available on the data system.

POX11 I is not working. I didn't investigate what was wrong. Please make sure when you come to need POX11.

The orthogonalities of POY11 and POP55 were measured and already adjusted. The results are below:

POY11

ABS = 0.973633 

PHASE = 92.086483 [deg]

ezcawrite C1:LSC-POY11_Q_GAIN 1.027081 && ezcawrite C1:LSC-POY11_PHASE_D 92.086483

POP55

ABS =  1.02680579

PHASE =  88.5246 [deg]

ezcawrite C1:LSC-POP55_Q_GAIN 0.973894 && ezcawrite C1:LSC-POP55_PHASE_D 88.524609

 

 


  5440   Fri Sep 16 21:26:12 2011 KeikoUpdateLSC3f demodulation board check

The demodulation phases and gains for the all existing channels, AS11, REFL11,REFL55, REFL165, and REFL33, were adjusted by the command "ezcawrite" commands. 

Scripts are: 

REFL165 ezcawrite C1:LSC-REFL165_Q_GAIN 0.934340 && ezcawrite C1:LSC-REFL165_PHASE_D -81.802479

REFL33

ezcawrite C1:LSC-REFL33_Q_GAIN 0.984244 && ezcawrite C1:LSC-REFL33_PHASE_D -89.618

REFL11

ezcawrite C1:LSC-REFL11_Q_GAIN 1.173418 && ezcawrite C1:LSC-REFL11_PHASE_D -442.882697

AS11
ezcawrite C1:LSC-AS11_Q_GAIN 0.975576 && ezcawrite C1:LSC-AS11_PHASE_D -93.12492

AS55

ezcawrite C1:LSC-AS55_Q_GAIN 0.999164 && ezcawrite C1:LSC-AS55_PHASE_D -89.300986
  5439   Fri Sep 16 17:46:13 2011 kiwamuUpdateSUSSome screens fixed

The bad medm screens have been fixed. There are no blank fields and all the links are correct.

Quote from #5409

I've found that a few of the screens still have Whited-Out fields due to naming changes (OL SUM and ALS-> TM OFFSET). I attach a screen shot of it.

The OL screens have the wrong SUM names and the IFO ALIGN screen is pointing to the wrong SUS screens.

 

  5438   Fri Sep 16 17:16:15 2011 JenneUpdateSUSInput matrix diagonalization: Fail!

[Jenne, Anamaria]

I put the new matricies in from the free swinging test for the: ITMX, ITMY, ETMX, ETMY, PRM, BS

Some of the optics damped okay, but ETMX and BS were not good at all.  ETMX was ringing up when I turned on the damping.  BS wasn't, but when I gave it a kick, it wouldn't damp.  No good.

I tried ITMY, and it was totally fine, with nice damping Qs of ~5.  So, I don't know what's going on. 

Anamaria is trying a new 4x4 matrix-inverter, so we can look at the inversion of just the face osems.  We'll see how it goes. 

Since things were crappy, I did a BURT restore, so things are as they were earlier this morning.

  5437   Fri Sep 16 17:09:07 2011 PaulUpdateSUSITMX oplev plan

 I just drew a basic picture of how the ITMX oplev path could be reworked to minimise the number of optics in the path. Only possible problem with this might be the turning mirror onto the ITMX getting in the way of the collimating lenses. Should be easy to solve though. Does anyone know if there is a ITMX pick off beam I should be careful to avoid?

Attachment 1: ITMX_oplev_plan.png
ITMX_oplev_plan.png
  5436   Fri Sep 16 16:34:54 2011 PaulUpdateSUSITMY SRM oplev telescope plan

I've calculated a suitable collimating telescope for the ITMY/SRM oplev laser, based on the specs for the soon-to-arrive 2mW laser (model 1122/P) available here: http://www.jdsu.com/ProductLiterature/hnlh1100_ds_cl_ae.pdf

Based on the fact that the 'beam size' value and 'divergence angle' value quoted don't match up, I am assuming that the beam radius value of 315um is _not_ the waist size value, but rather the beam size at the output coupler. From the divergence angle I calculated a 155um waist, (zR = 12cm). This gives the quoted beam size of about 316um at a distance of 8.5" away from the waist. This makes me think that the output coupler is curved and the waist is at the back of the laser, or at least 8.5" from the output coupler.

The collimating telescope gives a waist of size 1142um (zR=6.47m) at a distance of 1.427m away from the original laser waist, using the following lens combo:

 

L1 f=-0.15 @ 0.301m

L2 f=0.3 @ 0.409m

 

This should be fine to get a small enough spot size (1-2mm) on the QPDs.

 

Attachment 1: ITMY_SRM_telescope.png
ITMY_SRM_telescope.png
  5435   Fri Sep 16 16:29:05 2011 kiwamuUpdateSUSf2a filters on SRM

New f2a filters were installed on SRM.

The lock of DRMI should be more stable than last night.

Quote from #5417

Once the SRM oplev project settles down, I will adjust the f2a filters on SRM too.

 

Attachment 1: F2ASRM_Sep16.png
F2ASRM_Sep16.png
  5434   Fri Sep 16 16:07:28 2011 steveUpdateSAFETYvisitors safety training

Paul, Mirko and Katrin visiting grad students received the 40m basic safety training.

Attachment 1: P1080241.JPG
P1080241.JPG
  5433   Fri Sep 16 14:43:43 2011 SureshUpdateComputer Scripts / ProgramsChanged c1ioo model and restarted fb

I had to change the c1ioo model and restart the fb since the paths allowing us to select various signals to demodulate using the lockins were not correct. The signal selection vector was not flexible enough to permit us to select the signals to demod.

fb was restarted twice at following times.  The changes have been commited to the svn.

Fri Sep 16 13:35:47 PDT 2011

Fri Sep 16 14:36:21 PDT 2011


 

  5432   Fri Sep 16 14:03:53 2011 PaulUpdateSUSSRM oplev QPD noise measurement

 I checked the dark and bright noise of the SRM oplev QPD. The SRM QPD has a rather high dark level for SUM of 478 counts. The dark noise for the SRM QPD looked a little high in the plot against the bright noise (see first attachment), so I plotted the dark noise with the ITMY QPD dark noise (see second attachment). It seems that the SRM QPD has a much higher dark noise level than the ITMY! In case anyone is wondering, to make these traces I record the data from the pitch and yaw test points, then multiply by the SUM (to correct for the fact that the test point signal has already been divided by SUM). I will check the individual quadrants of the SRM QPD to see if one in particular is very noisy. If so, we/I should probably fix it.

Attachment 1: SRM_oplev_dark_noise_vs_bright_noise.pdf
SRM_oplev_dark_noise_vs_bright_noise.pdf
Attachment 2: SRM_ITMY_QPD_dark_noise_comparison.pdf
SRM_ITMY_QPD_dark_noise_comparison.pdf
  5431   Fri Sep 16 11:15:12 2011 KojiUpdateIOOPZT1 situation

[Koji Kiwamu]

- We have checked the situation of the broken Piezo Jenna PZT (called PZT1)

- Tested PZT1 by applying a dc voltage on the cables. Found that pitch and yaw reasonably moving and the motions are well separated each other.
  The pitch requires +20V to set the IPPOS spot on the QPD center.

- Made a high-voltage (actually middle voltage) amp to convert +/-10V EPICS control signal into -5 to +30V PZTout. It is working on the prototype board and will be put into the actual setup soon.


Details:

- The Piezo Jenna driver box has 4 modules. From the left-hand side, the HV module, Yaw controller, Pitch controller, and Ben abbot's connector converter.

- We checked the voltage on Ben's converter board. (Photo1)
  It turned out that the red cable is the one have the driving voltage while the others stays zero.

- We hooked a 30V DC power supply between the red cable and the shield which is actually connected to the board ground.

- Applying +/-10V, we confirmed the strain gauge is reacting. If we actuated the pitch cable, we only saw the pitch strain gauge reacted. Same situation for yaw too.

- Kiwamu went to IPPOS QPD to see the spot position, while I change the voltage. We found that applying +20V to the pitch cable aligns the spot on the QPD center.

------------------------

- I started to make a small amplifier boards which converts +/-10V EPICS signals into -5V to +30V PZT outs.

- The OPAMP is OPA452 which can deal with the supply voltages upto +/-40V. We will supply +/-30V. The noninerting amp has the gain of +2.

- It uses a 15V zener diode to produce -15V reference voltage from -30V. This results the output voltage swing from -5V to +35V.
The actual maximum output is +30V because of the supply voltage.

- On the circut test bench, I have applied +/-5V sinusoidal to the input and successfully obtained +5V to +25V swing.

- The board will be put on Ben's board today.

Attachment 1: P9151574.JPG
P9151574.JPG
Attachment 2: P9161576.JPG
P9161576.JPG
Attachment 3: P9161577.JPG
P9161577.JPG
  5430   Fri Sep 16 03:22:11 2011 AnamariaUpdateLSCMore Refl PDs Work and Attempt at DRMI

Kiwamu, Keiko, Anamaria

I started today with a different input beam, so I had to realign the REFL path again. Then we measured the RF signal out of the 4 REFL PDs and found them to be too low. We increased the power to around 10mA for each diode, and we can see the right modulation frequency on each diode, though REFL165 is way too weak so we might need an RF amplifier on it. We will measure demod board noise tomorrow.

We had an issue with REFL165 not giving the right DC level, low by a factor of 10, even though it was receiving the same optical power as the others. We fifteen-checked clipping and alignment, then pulled it out and measured it on the test stand - found it to be ok. So I uplugged its power cable at the rack and connected it to the AS165 slot. Problem sloved. Not sure what was wrong with the other power slot.

Then we found REFL55 to be clipping on its black glass, we fixed that. But the REFL55 DC power still changes a lot with seemingly not huge motions of the PRM. We'll investigate more tomorrow.

We added a lens in the path to REFL165 because unlike the others it is a 1mm diode. All diodes have about half a turn to a full turn flatness of maximum (on tiny steering mirror).

We set the whitening gain on all four diodes to 21 db.

Not sure if we should set the power to be different on these diodes since their sensitivity is different to RF, and now REFL11 sees huge signal.

We continued the DRMI locking attempt and brought in the SRC, using AS55I to control it. It kind of works/stays locked. We did manage to get MICH and PRC better controlled than last night, but with SRC in the mix, something is wrong. We have to redo f2a filters on SRM and hopefully things will be better after Jenne's suspension work tomorrow. Oplevs not optimized yet either.

We intend to realign POY beam path so we can monitor power in cavities.

  5429   Fri Sep 16 00:08:30 2011 PaulUpdateSUSITMY Oplev QPD dark and bright noise spectra

 I tried again at plotting the ITMY_QPD noise spectra in for dark and bright operation. Before we had the strange situation where the dark noise seemed higher, but Kiwamu noticed this was caused by dividing by the SUM before the testpoint I was looking at. This time I tried just multiplying by the measured SUM for bright and dark to normalise the spectra against each other. The results looks more reasonable now, the dark noise is lower than the bright noise for a start! However, the dark noise spectrum now doesn't look the same as the one I showed in my previous post.

Attachment 1: ITMY_oplev_dark_noise_vs_bright_noise.pdf
ITMY_oplev_dark_noise_vs_bright_noise.pdf
  5428   Thu Sep 15 22:31:44 2011 ManuelUpdateSUSSummary screen

I changed some colors on the Summary of Suspension Sensor  using my italian creativity.

I wrote a script in Python to change the thresholds for the "alarm mode" of the screen.

The script takes a GPS-format start time as the 1st argument and a duration time as the second argument.

For every channel shown in the screen, it compute the mean value during this time.

The 3rd argument is the ratio between the mean and the LOW threshold. The 4th argument is the ratio between the mean and the LOLO threshold.

Then it sets the thresholds simmetrycally for HIGH and HIHI threshold.

It does that for all channels skipping the Gains and the Off Sets because this data are not stored.

For example is ratio are 0.9 and 0.7 and the mean is 10, thresholds will be LOLO=7, LOW=9, HIGH=11, HIHI=13.

You can run the script on pianosa writing on a terminal '/opt/rtcds/caltech/c1/scripts/SUS/set_thresholds.py' and the arguments.

I already run my program with those arguments: 1000123215 600 0.9 0.7

The time is of this morning at 5:00 for 10 minutes

 

This is the help I wrote

HELP: This program set the thresholds for the "alarm mode" of the C1SUS_SUMMARY.adl medm screen.

 Written by Manuel Marchio`, visiting student from University of Pisa - INFN for the 2011 summer at Ligo-Caltech. Thrusday, 15th September 2011.

The 1st argument is the time in gps format when you want to START the mean

The 2nd argument is the DURATION

The 3rd argument is the ratio of the LOW and the HIGH thresholds. It must be in the range [0,1]

The 4th argument is the ratio of the LOLO and the HIHI thresholds. It must be in the range [0,1]

Example: path/set_thresholds.py 1000123215 600 0.9 0.7

and if the the mean is 10, thresholds will be set as LOLO=7, LOW=9, HIGH=11, HIHI=13

 

Attachment 1: sussum.png
sussum.png
  5427   Thu Sep 15 22:26:32 2011 PaulUpdateSUSITMY Oplev QPD dark noise PSD

 I took a dark noise measurement for the ITMY QPD, for comparison with measurements of the oplev noise later on. Initially I was plotting the data from test points after multiplication by the oplev matrix (i.e. the OLPIT_IN1 / OLYAW_IN1), but found that the dark noise level seemed higher than the bright noise level (!?). Kiwamu realised that this is because at that test point the data is already divided by QPD SUM, thus making the dark noise level appear to be greater than the bright level, since QPD SUM is much smaller for the dark measurements. The way around this was to record the direct signals from each quadrant before the division. I took a power spectrum of the dark noise from each quadrant, then added them in quadrature, then divided by QPD SUM at the end to get an uncalibrated PSD. Next I will convert these into the equivalent for pitch and yaw noise spectra. To calibrate the plots in radians per root Hz requires some specific knowledge of the oplev path, so I won't do this until I have adjusted the path.

Attachment 1: ITM_dark_QPD_PSD.pdf
ITM_dark_QPD_PSD.pdf
  5426   Thu Sep 15 21:56:01 2011 MirkoUpdateCDSc1oaf check, possible shmem problem

After Jamie installed the c1oaf model ( entry 5424 ) I went and checked the intermodel communication.

Remember the config is:

c1lsc ->SHMEM-> c1oaf
c1oaf ->SHMEM-> c1lsc
c1pem ->SHMEM-> c1rfm ->PCIE-> c1oaf

I checked at least one of every communications type.

-All signals reach their destinations.
-c1lsc_to_c1oaf_via_shmem is more noisy adding noise to the signal. lsc runs at 16kHz and oaf at 2kHz but that should actually smooth things out.

c1lsc_to_c1oaf_via_shmem.png

 

Attachment 1: c1lsc_to_c1oaf_via_shmem.png
c1lsc_to_c1oaf_via_shmem.png
Attachment 2: c1oaf_to_c1lsc_via_shmem_fixed_sine_inj_at_100Hz.png
c1oaf_to_c1lsc_via_shmem_fixed_sine_inj_at_100Hz.png
Attachment 3: c1oaf_to_c1lsc_via_shmem_white_noise_inj.png
c1oaf_to_c1lsc_via_shmem_white_noise_inj.png
Attachment 4: c1pem_to_c1oaf_via_rfm.png
c1pem_to_c1oaf_via_rfm.png
  5425   Thu Sep 15 20:30:52 2011 MirkoUpdateLSCPlan for long term optical spec. analyzer setup

Just to give some heads up on how the setup on the PSL table does / will look. We start out with one of the two reflections of a window. Power about 2mW.

DSC_3430b.jpg

  5424   Thu Sep 15 20:16:15 2011 jamieUpdateCDSNew c1oaf model installed and running

[Jamie, Jenne, Mirko]

New c1oaf model installed

We have installed the new c1oaf (online adaptive feed-forward) model.  This model is now running on c1lsc.  It's not really doing anything at the moment, but we wanted to get the model running, with all of it's interconnections to the other models.

c1oaf has interconnections to both c1lsc and c1pem via the following routes:

c1lsc ->SHMEM-> c1oaf
c1oaf ->SHMEM-> c1lsc
c1pem ->SHMEM-> c1rfm ->PCIE-> c1oaf

Therefore c1lsc, c1pem, and c1rfm also had to be modified to receive/send the relevant signals.

As always, when adding PCIx senders and receivers, we had to compile all the models multiple times in succession so that the /opt/rtcds/caltech/c1/chans/ipc/C1.ipc would be properly populated with the channel IPC info.

Issues:

There were a couple of issues that came up when we installed and re/started the models:

c1oaf not being registered by frame builder

When the c1oaf model was started, it had no C1:DAQ-FB0_C1OAF_STATUS channel, as it's supposed to.  In the daqd log (/opt/rtcds/caltech/c1/target/fb/logs/daqd.log.19901) I found the following:

Unable to find GDS node 22 system c1oaf in INI files

It turns out this channel is actually created by the frame builder, and it could not find the channel definition file for the new model, so it was failing to create the channels for it.  The frame builder "master" file (/opt/rtcds/caltech/c1/target/fb/master) needs to list the c1oaf daq ini files:

/opt/rtcds/caltech/c1/chans/daq/C1OAF.ini
/opt/rtcds/caltech/c1/target/gds/param/tpchn_c1oaf.par

These were added, and the framebuilder was restarted.  After which the C1:DAQ-FB0_C1OAF_STATUS appeared correctly.

SHMEM errors on c1lsc and c1oaf

This turned out to be because of an oversight in how we wired up the skeleton c1oaf model.  For the moment the c1oaf model has only the PCIx sends and receives.  I had therefore grounded the inputs to the SHMEM parts that were meant to send signals to C1LSC.  However, this made the RCG think that these SHMEM parts were actually receivers, since it's the grounding of the inputs to these parts that actually tells the RCG that the part is a receiver.  I fixed this by adding a filter module to the input of all the senders.

Once this was all fixed, the models were recompiled, installed, and restarted, and everything came up fine.

All model changes were of course committed to the cds_user_apps svn as well.

  5423   Thu Sep 15 18:31:27 2011 PaulUpdateSUSITMY and SRM Oplev current status - comparison with ITMY

Quote:

Just to find out where we are currently, I plotted the ITMY and SRM oplev spectra along with the ETMY oplev spectra. ETMY seems to be very good, so comparing with this seemed useful, so we know how much we have to improve by. The SRM power spectrum appears to be around 2 orders of magnitude higher than ETMY over pretty much the whole measurement band. The ITMY power spectrum is not so bad as the SRM above about 60Hz. Next thing to do is to check the dark noise level for the ITMY and SRM QPDs.

 The title of this post should of course have been " ... - comparison with ETMY" not " ... - comparison with ITMY"

  5422   Thu Sep 15 18:24:54 2011 PaulUpdateSUSITMY and SRM Oplev current status - comparison with ITMY

Just to find out where we are currently, I plotted the ITMY and SRM oplev spectra along with the ETMY oplev spectra. ETMY seems to be very good, so comparing with this seemed useful, so we know how much we have to improve by. The SRM power spectrum appears to be around 2 orders of magnitude higher than ETMY over pretty much the whole measurement band. The ITMY power spectrum is not so bad as the SRM above about 60Hz. Next thing to do is to check the dark noise level for the ITMY and SRM QPDs.

Attachment 1: oplev_spectra_comparison.pdf
oplev_spectra_comparison.pdf
  5421   Thu Sep 15 18:12:21 2011 JenneUpdateSUSfree swinging test in vacuum condition

Quote:

All the optcs were excited

Sat Sep 10 02:14:11 PDT 2011
999681266

 

 

Optic The Plot Input Matrix BADness
ITMX  ITMX.png       pit     yaw     pos     side    butt
UL    0.601   0.680   1.260  -1.009   0.223 
UR    0.769  -1.254  -0.175  -0.179   0.581 
LR   -1.231   0.065   0.566  -0.480   0.252 
LL   -1.399   2.000   2.000  -1.310  -2.944 
SD   -0.580   0.868   2.451   1.000  -1.597 

 
7.95029
ITMY  ITMY.png       pit     yaw     pos     side    butt
UL    1.067   0.485   1.145  -0.195   0.929 
UR    0.548  -1.515   0.949  -0.142  -1.059 
LR   -1.452  -0.478   0.855  -0.101   1.051 
LL   -0.933   1.522   1.051  -0.153  -0.962 
SD   -0.530   0.903   2.115   1.000   0.142 
3.93939
ETMX ETMX.png       pit     yaw     pos     side    butt
UL    0.842   1.547   1.588  -0.018   1.026 
UR    0.126  -0.453   1.843   0.499  -1.173 
LR   -1.874  -0.428   0.412   0.511   0.934 
LL   -1.158   1.572   0.157  -0.006  -0.867 
SD    1.834   3.513  -0.763   1.000  -0.133
5.39825
ETMY ETMY.png       pit     yaw     pos     side    butt
UL   -0.344   1.280   1.425  -0.024   0.903 
UR    1.038  -0.720   1.484  -0.056  -1.161 
LR   -0.618  -1.445   0.575  -0.040   0.753 
LL   -2.000   0.555   0.516  -0.007  -1.184 
SD   -0.047  -0.038   0.986   1.000   0.083 
4.15747
BS  BS.png       pit     yaw     pos     side    butt
UL    1.549   0.655   0.393   0.263   0.997 
UR    0.192  -1.345   1.701  -0.063  -0.949 
LR   -1.808  -0.206   1.607  -0.085   0.952 
LL   -0.451   1.794   0.299   0.241  -1.101 
SD    0.724   0.293  -3.454   1.000   0.037 
5.66432
PRM  PRM.png       pit     yaw     pos     side    butt
UL    0.697   1.427   1.782  -0.337   0.934 
UR    1.294  -0.573   0.660  -0.068  -0.943 
LR   -0.706  -1.027   0.218   0.016   0.867 
LL   -1.303   0.973   1.340  -0.254  -1.257 
SD    0.369  -0.448  -0.496   1.000   0.456 
5.1026
SRM   Can't invert....need to fix the peak-finding.  
MC1  MC1.png       pit     yaw     pos     side    butt
UL    0.872   0.986   0.160   0.054   0.000 
UR    0.176  -0.752   0.917   0.018   0.000 
LR   -1.824  -2.000   1.840   0.002   3.999 
LL   -1.128  -0.262   1.083   0.038  -0.000 
SD    0.041   0.036  -0.193   1.000  -0.001 
5.31462
MC2  MC2.png       pit     yaw     pos     side    butt
UL    1.042   0.767   0.980   0.131   0.928 
UR    0.577  -1.233   1.076  -0.134  -0.905 
LR   -1.423  -0.640   1.020  -0.146   1.050 
LL   -0.958   1.360   0.924   0.120  -1.117 
SD   -0.073  -0.164  -0.702   1.000  -0.056 
4.07827
MC3  MC3.png       pit     yaw     pos     side    butt
UL    1.595   0.363   1.152   0.166   1.107 
UR    0.025  -1.629   1.135   0.197  -0.994 
LR   -1.975   0.008   0.848   0.105   0.904 
LL   -0.405   2.000   0.865   0.074  -0.995 
SD   -0.433   0.400  -1.624   1.000   0.022 
3.64881

 

  5420   Thu Sep 15 17:12:29 2011 MirkoUpdateLSCRF modulation depth measurement

[Mirko, Kiwamu]

Put up a temp. setup on the laser table to measure the RF modulation depths using the optical spectrum analyzer. First with a pickoff beam with about 2mW => SNR of 8 of 1 peak per FSR.

Then with a beam with about 100mW. Much better SNR on the single peak but still no sidebands visible. Modematching not too good in either case. Shouldn't matter.

  5419   Thu Sep 15 17:00:10 2011 Paul and SteveUpdateSUSNew ITMY and SRM oplev plan

 We have made a new plan for the ITMY and SRM oplev optical path which uses as few optics as possible. This should help to reduce coupling from vibrations of optics in the oplev path back into the GW channel. To get enough room for the turning mirror into the SRM it might be necessary to move the POY optics a bit nearer to the tank. 

Attachment 1: oplev_plan1.png
oplev_plan1.png
  5418   Thu Sep 15 16:45:59 2011 PaulUpdateSUSITMY and SRM Oplev status

Today I worked on getting the ITMY and SRM oplevs back in working order. I aligned the SRM path back onto the QPD. I put excitations on the ITMY and SRM in pitch and yaw and observed the beam at the QPDs to check for clipping. They looked clean from clipping.

 
Measurements of the beam power at various points:
 
Straight after the laser - 7.54mW
After the BS in the SRM path - 1.59mW
After the BS in the ITMY path - 3.24mW
Incident on the SRM QPD - 0.03mW
Incident on the ITMY QPD - 0.25mW
 
Counts registered from the QPD sum channels:
 
SRM QPD SUM dark count - 1140
SRM QPD SUM bright count - 3250
 
ITMY QPD SUM dark count - 150
ITMY QDP SUM bright count - 12680
 
The power incident on the SRM QPD seems very low with respect to the ITMY QPD. Is the SRM mirror coating not very reflective for the He-Ne laser?There are some back reflections from lenses, which we should be careful of to avoid scattering.
  5417   Thu Sep 15 15:11:38 2011 kiwamuUpdateSUSf2a filters on BS and PRM

The f2a filters were newly designed and installed on BS and PRM.

So the lock of PRMI will be more stable .

Once the SRM oplev project settles down, I will adjust the f2a filters on SRM too.

Attachment 1: PRMf2a.png
PRMf2a.png
Attachment 2: BS_f2a.png
BS_f2a.png
  5416   Thu Sep 15 11:37:24 2011 SureshUpdateComputer Scripts / Programsfb restarted at Thu Sep 15 11:30:30 PDT 2011

I changed a filter bank name (C1IOO-WFS1_PIT) in c1ioo model reverting it to its earlier name.  Had to restart c1ioo model and the fb

  5415   Thu Sep 15 07:28:08 2011 steveUpdateSUSPRM damping restored

Quote:

The PRM damping was restored at side sensor var 1050

 The PRM sus damping restored.

  5414   Thu Sep 15 02:18:19 2011 AnamariaUpdateLSCMICH locked and attempt to lock PRCL

Kiwamu, Keiko, Anamaria

 

We were able to lock PRC using REFL11I after improving the MICH dark fringe a bit (moving BS) and rotating AS55 and REFL11 such that the signal was maximized in the phases we were using. The dark port is not so dark... but the lock is stable.

I had finished the whole REFL path alignment, but I didn't have a good input beam reference at the time, which is why we had to realign the PDs and the camera. We only had strength to realign 11 and 55. Otherwise, we just need to tweak and center beam on 33 and 165, figure out what's up with 55 and be done with the AP table mods. I hope.

 

Quote:

 Anamaria, Keiko

- We aligned MICH and were successfully locked MICH using AS55Q. The other mirrors were misaligned so that the other degrees of freedom didn't exist. AS55 was fed back to BS. The f2a filters on BS suspension were required to lock, because the pos feedback was unbalanced to angle degrees of freedom.

- We tried to lock PRCL next, however, because we aligned the MICH and the REFL beam paths were changed, REFL PDs didn't have the light anymore. The REFL paths were modified now, so we will try the PRCL locking next.

- We couldn't confirm REFL55 signals although we alined the REFL paths to REFL55 PD.

 

  5413   Thu Sep 15 01:17:10 2011 KeikoUpdateLSCMICH locked and attempt to lock PRCL

 Anamaria, Keiko

- We aligned MICH and were successfully locked MICH using AS55Q. The other mirrors were misaligned so that the other degrees of freedom didn't exist. AS55 was fed back to BS. The f2a filters on BS suspension were required to lock, because the pos feedback was unbalanced to angle degrees of freedom.

- We tried to lock PRCL next, however, because we aligned the MICH and the REFL beam paths were changed, REFL PDs didn't have the light anymore. The REFL paths were modified now, so we will try the PRCL locking next.

- We couldn't confirm REFL55 signals although we alined the REFL paths to REFL55 PD.

  5412   Thu Sep 15 01:06:20 2011 KeikoUpdateLSC3f demodulation board check

In addition to REFL 33 ans 165, I checked the orthogonality for the other existing three channels.

 

AS11

ABS = 1.025035  PHASE = -93.124929 [deg]

REFL11]

ABS = 0.920984  PHASE = -88.824691 [deg]

REFL55

ABS = 1.029985 , PHASE = -90.901123 [deg]

- - - - - - - - - - - - - - - - - -

 

The demodulated signal was set as 50 Hz (for example LO 11MHz and RF 11MHz+50Hz from function generators.) These AS11, REFL11, REL55, REFL33m REFL165 are the current available channels in terms of the connection to the data system from the demodulation board. I am going to estimate the error next.

Quote:

REFL165

ABS = 1.070274, PHASE = -81.802479 [deg]

- - - - - - - - - - - - - - - - - -

REFL33

ABS = 1.016008 PHASE = -89.618724 [deg]

 

  5411   Wed Sep 14 22:07:41 2011 ranaUpdateSUSITM Oplevs are broken

I went to see what was wrong with the ITMs and found that people have been working on them and have left them in a broken state with no elog entry.

This is sad and unacceptable.

Whoever is working on these should post into the elog what the Oplev layout plan is, have someone check it, and ONLY THEN get to work on it.

The layout as it looks tonight is too complicated. With too many optics we will not have a low noise optical lever setup. The new layout should use a bare minimum number of optics and only use very stable mounts.

mm.jpg

  5410   Wed Sep 14 21:18:45 2011 kiwamuUpdateLSCY arm locking prep

Although we did some of the Input Matrix diagonalization, we have not yet actually used this knowledge. As a result all of the optics are shaking all over the place.

Sunshine Task: Set the input matrices to their calculated values and then adjust the OSEM damping gains for all optics so as to get a Q ~ 5.

  5409   Wed Sep 14 20:30:36 2011 ranaUpdateSUSSome screens are still bad

I've found that a few of the screens still have Whited-Out fields due to naming changes (OL SUM and ALS-> TM OFFSET). I attach a screen shot of it.

The OL screens have the wrong SUM names and the IFO ALIGN screen is pointing to the wrong SUS screens.

Untitled.png

  5408   Wed Sep 14 20:04:05 2011 jamieUpdateCDSUpdate to frame builder wiper.pl script for GPS 1000000000

I have updated the wiper.pl script (/opt/rtcds/caltech/c1/target/fb/wiper.pl) that runs on the framebuilder (in crontab) to delete old frames in case of file system overloading.  The point of this script is to keep the file system from overloading by deleting the oldest frames.  As it was, it was not properly sorting numbers which would have caused it to delete post-GPS 1000000000 frames first.  This issue was identified at LHO, and below is the patch that I applied to the script.


--- wiper.pl.orig  2011-04-11 13:54:40.000000000 -0700
+++ wiper.pl       2011-09-14 19:48:36.000000000 -0700
@@ -1,5 +1,7 @@
 #!/usr/bin/perl
 
+use File::Basename;
+
 print "\n" .  `date` . "\n";
 # Dry run, do not delete anything
 $dry_run = 1;
@@ -126,14 +128,23 @@
 
 if ($du{$minute_trend_frames_dir} > $minute_frames_keep) { $do_min = 1; };
 
+# sort files by GPS time split into prefixL-T-GPS-sec.gwf
+# numerically sort on 3rd field
+sub byGPSTime {
+    my $c = basename $a;
+    $c =~ s/\D+(\d+)\D+(\d+)\D+/$1/g;
+    my $d = basename $b;
+    $d =~ s/\D+(\d+)\D+(\d+)\D+/$1/g;
+    $c <=> $d;
+}
+
 # Delete frame files in $dir to free $ktofree Kbytes of space
 # This one reads file names in $dir/*/*.gwf sorts them by file names
 # and progressively deletes them up to $ktofree limit
 sub delete_frames {
        ($dir, $ktofree) = @_;
        # Read file names; Could this be inefficient?
-       @a= <$dir/*/*.gwf>;
-       sort @a;
+       @a = sort byGPSTime <$dir/*/*.gwf>;
        $dacc = 0; # How many kilobytes we deleted
        $fnum = @a;
        $dnum = 0;
@@ -145,6 +156,7 @@
          if ($dacc >= $ktofree)  { last; }
          $dnum ++;
          # Delete $file here
+         print "- " . $file . "\n";
          if (!$dry_run) {     
            unlink($file);
          }

  5407   Wed Sep 14 15:05:52 2011 steveUpdatePEMair cond maintenance

Jeff has changed our AC filters inside the lab this morning. Now he is checking on the main filters at CES.  He will finish the roof units tomorrow.

Met One #1 counter is on the top of IOO chamber.  It is measuring 1 and 0.5 micron size particles.  One year of lab condition plot below.

 

Attachment 1: ac1y.jpg
ac1y.jpg
  5406   Wed Sep 14 15:00:00 2011 kiwamuUpdateLSCmeasurement of spot position on Y arm

Forgot to attach a picture of the ITMY's face camera when it was locked.

The horizontal position of the spot looks good, but the vertical position is apparently too low, which agrees with the A2L result.

DSC_3408_small.jpg

Quote from #5400
  amount of off-centering [mm]
ETMY_PIT  0.0032
ETMY_YAW  -0.13
ITMY_PIT  -12
ETMY_YAW  0.91

 

 

  5405   Wed Sep 14 14:06:44 2011 kiwamuUpdateGeneralminutes of 40m meeting : plans
+ Optical lever (Steve/Paul)
+ LSC demod board (Keiko)
+ DRMI locking (Keiko / Anamaria)
+ Modulation depth (Mirko)
+ ALS (Katrin)
+ MC WFS (Suresh)
+ OAF (Jenne/Mirko)
+ MC_F acquisition (Jenne/Mirko)
+ SUS, free swinging (Jenne)
+ CDS (Jamie)
+ vent plan (Jamie)
+ EOM (Kiwamu)
+ PZT-HV (Kiwamu/Koji)
  5404   Wed Sep 14 12:01:05 2011 ranaUpdateAdaptive FilteringModifications to LSC, RFM models, added OAF model

For the acquisition of the MC_F channel, I suggest taking the FAST_MON BNC output from the blue FSS interface card in the Eurocard crate in the PSL rack. This can then be piped into the 2-pin LEMO plug (Ch. 1) of the Generic Pentek DAQ card which used to acquire the MC_L signal from the MC Servo Board.

  5403   Wed Sep 14 07:51:20 2011 steveUpdateSUSPRM damping restored

The PRM damping was restored at side sensor var 1050

  5402   Wed Sep 14 01:21:17 2011 JenneUpdateAdaptive FilteringModifications to LSC, RFM models, added OAF model

[Jenne, Mirko, with supervision from Jamie]

We are starting to create the new OAF model, so that it works with the new CDS system. 

I created (and did an "svn add" for) a new c1oaf.mdl, in the same place as the current c1lsc.mdl . Since the oaf will kind of be working with ISC things, I decided to put it in that folder.  So far this new OAF model just has SHMEM/PCIE memory sharing things to get info from the LSC and PEM models.  The OAF model has dcuid=22 (the same as in the old system), and lives on the LSC machine with specific_cpu=4.  This is the CPU number that Yoichi was going to use, but he ended up putting his FF stuff directly into the LSC model for delay reasons.

I modified the c1rfm.mdl to take seismometer and accelerometer info from the PEM model, and give it to the "rfm" via shmem, and then using PCIE (dolphin) to get the channel to the OAF model.

I modified the c1lsc model to have shmem outputs that go from the degrees of freedom to the OAF, and shmem inputs from the OAF's output to sum into the DoFs, just like Yoichi's FF stuff.  I also removed the old OAF_OUT, because it would only allow me to select one DoF at a time, and I will eventually want the ability to do multiple amounts of OAFing at the same time.  Hopefully.

All of the above changes have been svn'ed with nice log messages into the cds svn.

I have not yet modified the PEM model to give the seis/acc information to the RFM model.

I will need to acquire the PSL's PZT input as a representation of the mode cleaner's length if I want to apply the OAF to the MC to recreate past work.

  5401   Wed Sep 14 01:19:20 2011 AnamariaConfigurationLSC3f PD Install in Progress

I have reconfigured the refl beam path on the AP table to include REFL33 and REFL165. Would be done if we hadn't prepared P BSs instead of S, which required some serious digging to find two others. And if someone hadn't stolen our two 3m SMA cables that Keiko and I made on our previous visit and I had left with the 3f PDs. I don't expect them to reappear but if they do, it would be grand.

Note: Refl beam from ifo looks a bit high, ~1cm on the lens 20'' from output port. Not sure what that means about ifo alignment change, I've left it as is. When we know we have a good alignment, we should be able to easily realign the beam path if necessary. If it remains the same, we might want to change the lens height.

Done:

1) REFL11 and REFL55 are now hooked up and aligned in a low power beam. (I set the power as low as I could by eye to not risk burning the PDs during alignment)

2) The required BSs and REFL33 and REFL165 are in place, powered.

3) I have set them in a configuration such that the beam is the same distance from the main beam, to adjust beam size easily for all 4.

4) Camera has been moved from main beam to behind a steering mirror, ND filters removed, centered on camera.

To Do:

1) Find one more longish SMA cable.

2) Align beam on REFL33 and REFL165.

3) Check beam size carefully. (I get a plateau on the scope, and I can "hide" the beam on the PD, but it could be better. The path has become longer by ~5-8inches.)

4) Adjust power.

5) Redo layout diagram, post in wiki.

ELOG V3.1.3-